Determine the nature of the solutions 4x^2-9x=0
Answer by stanbon(75887)

Determine the nature of the solutions 4x^2-9x=0
---------------
a = 4 ; b = -9 ; c = 0
---
Discriminant = b^2-4ac
= 81 - 4*4*0
= 81
----------------
Since discriminant > 0, the solutions are
unequal and Real.
Cheers,
Stan H.